OpenStudy (anonymous):

solve each system of equations by the substitution method 2x-8y=7 3x+y=-2

OpenStudy (anonymous):

y=-2-3x 2x-8(-2-3x)=7 2x+16+24x=7 26x=-9 x=-9/26 y=.......

OpenStudy (anonymous):

Thank You! Can you help me with my other math that i am going to post

OpenStudy (anonymous):

solve each system by substitution be sure to give yhour solution as an ordered pair x-y=2 x+y=6